# Enhancing Outlier Knowledge for Few-Shot Out-of-Distribution Detection with Extensible Local Prompts

### v1

@@ 4.2 Local prompt enhanced regional regularization
Once augmented inputs are decided, corresponding local features are then used to optimize local prompts to identify ID and OOD samples from the perspective of regional enhancement. We aim at utilizing the local features extracted from image encoder in a fine way to (1) learn regional-related prompts that better characterize local similarity for both ID and OOD regions; (2) enhance detection ability with refined local features. Specifically, we design local prompts with learnable context words $\bm{t}_{\mathrm{in}}$ to represent local textual information for each known class. Moreover, on account of the existence of OOD regions, we build a few local negative prompts $\bm{t}_{\mathrm{neg}}$ to handle possible local outliers.
@@
Contrastive regularization
We propose regional contrastive regularization to enhance local textual prompts for better local information utilization. It is composed of local loss and local negative loss. Given a positive sample of training image, local loss is employed to learn regional information of known classes:
$\mathcal{L}_{\mathrm{pos}}(\bm{x},y)=-\mathrm{log}\frac{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\}}{\sum_{\mathrm{in}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{in}})/T)\}}+\sum_{\mathrm{neg}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}}})/T)\}},$ (4)
@@
$\mathcal{L}_{\mathrm{neg}}(\tilde{\bm{x}})=-\mathrm{log}\frac{\sum_{\mathrm{neg}}\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}})/T)\}}{\sum_{\mathrm{in}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{in}})/T)\}}+\sum_{\mathrm{neg}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}})/T)}\}}.$ (6)
@@
Diversity regularization
As local negative prompts are randomly initialized, more regularization has to be imposed on them to ensure their diversity. Therefore, in addition to the proposed contrastive regularization, we apply a diversity regularization on local negative prompts, as follows:
$\mathcal{L}_{\mathrm{reg}}=\frac{1}{N_{\mathrm{neg}}^{2}}\sum_{1\leq i,j\leq N_{\mathrm{neg}}}\mathrm{sim}(\bm{t}_{i},\bm{t}_{j}),$ (7)
@@ 4.3 Regional OOD score
$S_{\mathrm{R-MCM}}(\bm{x})=S_{\mathrm{MCM}}(\bm{x})+\mathcal{T}_{k}^{\mathrm{mean}}\{\frac{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)}{\sum_{j=1}^{C+N_{\mathrm{neg}}}{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{l},\bm{t}_{j}})/T)}\},$ (9)
@@
$h_{i}^{\mathrm{g}}(\bm{x})=\rm sim(\bm{z}^{\mathrm{g}},\bm{t}_{i}),$ (10)
$h_{i}^{\mathrm{l}}(\bm{x})=\mathcal{T}_{k}^{\mathrm{mean}}\{\rm exp(sim(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\},$ (11)
@@
MSP [8] 54.05 87.43 73.37 78.03 72.98 78.03 68.85 79.06 67.31 80.64
Energy [47] 29.75 94.68 53.18 87.33 56.40 85.60 51.35 88.00 47.67 88.90
KNN [48] 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS [4] 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
Textual-OE [24] 29.61 94.74 57.12 87.34 66.82 83.71 79.29 77.76 58.21 85.88
@@
CoOp [31] 18.95 95.52 29.58 92.90 38.72 89.64 48.03 85.87 33.82 90.98
Ours 9.65 97.87 20.40 95.57 29.39 92.67 51.20 88.00 27.66 93.53
LoCoOp [18] 21.67 95.69 22.98 95.07 31.41 92.10 49.79 87.85 31.46 92.68
Ours+LoCoOp 12.81 97.29 19.34 95.85 27.53 92.97 45.51 89.99 26.29 94.03
@@
CoOp [31] 14.60 96.62 28.48 92.65 36.49 89.98 43.13 88.03 30.67 91.82
Ours 8.71 98.10 23.97 94.85 32.50 92.32 47.93 89.04 28.27 93.58
LoCoOp [18] 16.05 96.86 23.44 95.07 32.87 91.98 42.28 90.19 28.66 93.52
Ours+LoCoOp 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
@@
MCM [16] zero-shot 67.01
GL-MCM [17] 67.01
KNN [48] full-tuning 79.64
NPOS [4] 79.42
LoCoOp [18] 4-shot 69.50
Ours 70.52
LoCoOp [18] 16-shot 71.86
Ours 74.24
Table 3: Comparison results of near OOD detection
tasks. We report 16-shot results for few-shot methods.
Our method can achieve performance against other methods.
@@
MCM [16] 5.00 98.71 12.51 97.70 8.75 98.21
GL-MCM [17] 10.10 98.04 9.00 98.62 9.55 98.33
LoCoOp [18] 11.20 97.49 12.00 97.79 11.60 97.64
Ours 3.90 99.06 6.20 98.84 5.05 98.95
5 Experiments
5.1 Experimental setup
Datasets
Following existing works [16, 17, 18], large scale ImageNet-1K [49] along with a 100-category subset of it denoted as ImageNet-100 are used as ID dataset and OOD dataset is a combination of iNaturalist [50], SUN [51], Places [52] and Texture [53]. In addition, We also use two semantically similar subsets of ImageNet-1k, i.e., ImageNet-10 and ImageNet-20, to evaluate near OOD detection performance [16].
Implementation details
We use CLIP-Base/16 as the backbone. Image encoder, text encoder, and global prompts are frozen and only local prompts are learnable. We set one local prompt for each known category (1000 in total for ImageNet-1k). $\lambda_{\mathrm{neg}}$ and $\lambda_{\mathrm{reg}}$ are $5$ and $0.5$, respectively. As the stability of temperature $T$ has been verified by previous research [16, 17], we set $T$ to be 1 by default.
Evaluation metrics
We report the following metrics for evaluation: (1) the area under the receiver operating characteristic curve (AUROC); (2) false positive rate of OOD samples when true positive rate of ID samples is 95% (FPR95); (3) in-distribution data classification accuracy (ID accuracy).
5.2 Main results
ImageNet-1k as ID dataset
We report 4 and 16-shot results on four common OOD datasets using ImageNet-1k as ID dataset. It can be seen in Tab. 1 that tuning local prompts only using hand-crafted global prompts (Ours) can achieve competitive results, especially in datasets with challenging fine local textures like iNaturalist. Concretely, we get impressive progress in 4-shot tuning (outperforming by 3.80% on FPR95). In 16-shot setting, our approach gets competitive results as well. All the results above strongly showcase the potential of regional enhancement for OOD detection as an orthogonal direction to global prompt optimization methods.
Extensive to global prompts
Benefiting from the convenient integration with well-trained global prompts, we combine the advantage of trained global prompts from LoCoOp [18] without further design (Ours+LoCoOp) and once again get substantial improvements on all datasets. For example, we surpass the state-of-the-art method by a large margin (5.17% on FPR95 and 1.35% on AUROC), even against previous methods with 16 shots. Specifically, extending local prompts to global prompts is helpful to datasets with major global representations like Texture, verifying the compatibility and potential of local prompt optimization. It is notable that the extension process is seamless and can be integrated with any global optimization prompts, strongly showing the extensibility of our approach. We also provide density map in Fig. 3 to illustrate the superiority of our method.
ID accuracy
In addition to OOD performance, it is also meaningful to evaluate ID accuracy as it is expected to ensure classification accuracy and separate outliers at the same time. However, weighing too much about separating outliers may narrow the inter-class distance, thus harming accuracy. Consequently, we evaluate ID accuracy for ImageNet-1k and the results shown in Tab. 3 conclude that our method significantly surpasses previous methods (1.02% and 2.38%, respectively). It demonstrates that the proposed regional enhancement is also beneficial to improve OOD separation while increasing inter-class distance, which is of great significance for real application.
Table 4: Experiments on different datasets. ImageNet-100 is used as ID dataset. 4-shot results are reported for few-shot tuning.
@@ 5.2 Main results
MCM [16] 18.13 96.77 36.45 94.54 34.52 94.36 41.22 92.25 32.58 94.48
LoCoOp [18] 44.94 93.31 30.70 94.70 34.31 93.93 55.92 90.01 41.47 92.99
Ours 6.99 98.05 24.34 96.03 26.04 95.49 36.32 93.67 23.42 95.81
@@
Near OOD detection tasks
Near OOD task typically evaluates semantically similar categories and is relatively more challenging. Following existing work [16], we conduct experiments on ImageNet-10 vs. ImageNet-20 to verify the effectiveness of the proposed approach. Results in Tab. 3 indicate that our method exceeds previous methods by a large margin. It can be observed that previous few-shot method fails in near OOD tasks. We attribute it to the refinement of local prompts by local outlier knowledge, which greatly assists in distinguishing semantically similar categories. By contrast, relying solely on global prompts may confuse the distribution. Therefore, it underlines that our local outlier enhancement approach is especially effective in semantically similar near OOD tasks.
ImageNet-100 as ID dataset
We also conduct experiments using ImageNet-100 as ID dataset. We compare with other zero-shot and few-shot methods and results are reported in Tab. 4. The substantial performance improvements (average 9.16% reduction on FPR95 and 1.33% promotion on AUROC) strongly demonstrate the effectiveness and transferability of our approach.
5.3 Ablation Study
We conduct various and comprehensive ablation studies to investigate the effectiveness of each component. We train a 4-shot model to showcase the function of components and hyper-parameters unless otherwise stated.
@@ 5.3 Ablation Study
Table 7: Influence of different OOD score strategies. LNP stands for local negative prompts.
Method iNaturalist SUN Places Texture Average
[…]
MCM [16] 27.70 94.66 31.30 93.68 39.50 90.86 38.61 91.79 34.27 92.75
GL-MCM [17] 9.69 97.80 26.27 94.22 34.78 91.33 36.63 91.65 26.84 93.75
Ours(w/o LNP) 7.71 98.18 23.78 94.98 31.85 92.28 35.76 91.79 24.77 94.31
Ours(w LNP) 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
Effectiveness of loss components
We conduct experiments to analyse the effectiveness of each loss component and results in Tab. 5 imply that local prompts optimization (baseline) achieves comparable results with global prompts optimization methods, strongly demonstrating the effectiveness of our proposed regional enhancement strategy. Furthermore, taking hard negative samples together with local negative prompts $\bm{t}_{\mathrm{neg}}$ to optimize OOD-considered representation is beneficial to the distinction of outliers. Outcome without regularization loss is better in FPR95 but slightly inferior in AUROC. Considering overall performance and the fact that the optimization of $\bm{t}_{\mathrm{neg}}$ relies heavily on the initialization without regularization, we reserve regularization loss for stability and convergence.
Influence of negative augmentation strategy
We conduct experiments to analyse the influence of negative augmentation strategy. In Tab. 6, baseline does not employ augmentation. Cutout and random crop are used as augmentation, respectively. It is concluded that improper augmentation is unfavorable to detection ability. By contrast, random crop is helpful to enhance regional-related representation and distinguish outliers, which confirms our analysis above.
OOD score strategy
We compare our proposed OOD score with existing scores [16, 17] on 16-shots. We also remove local negative prompts in our OOD score to analyse their function. Results in Tab. 7 imply that our regional OOD score attains better performance than existing strategy, displaying the strength of regional enhancement. Moreover, the performance is further promoted when negative prompts are introduced. We attribute it to possible outlier leverage through negative prompts.

### v2

@@ 4.2 Local prompt enhanced regional regularization
Once augmented inputs are decided, corresponding local features are then used to optimize local prompts to identify ID and OOD samples from the perspective of regional enhancement. We aim at utilizing the local features extracted from image encoder in a fine way to (1) learn regional-related prompts that better characterize local similarity for both ID and OOD regions; (2) enhance detection ability with refined local features. Specifically, we design local prompts $\bm{t}$ with learnable context words to represent local textual information for each known class. Moreover, on account of the existence of OOD regions, we build a few local negative prompts $\bm{\hat{t}}$ to handle possible local outliers.
@@
Contrastive regularization. We propose regional contrastive regularization to enhance local textual prompts for better local information utilization. It is composed of local loss and local negative loss. Given a positive sample of training image, local loss is employed to learn regional information of known classes:
$\mathcal{L}_{\mathrm{pos}}(\bm{x},y)=-\mathrm{log}\frac{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{y})/T)\}}{\sum_{i=1}^{\mathrm{C}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\}}+\sum_{i=1}^{\mathrm{N_{neg}}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{\hat{t}}_{i})/T)\}}},$ (4)
@@
$\mathcal{L}_{\mathrm{neg}}(\tilde{\bm{x}})=-\mathrm{log}\frac{\sum_{i=1}^{\mathrm{N_{neg}}}\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{\hat{t}}_{i})/T)\}}{\sum_{i=1}^{\mathrm{C}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\}}+\sum_{i=1}^{\mathrm{N_{neg}}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{\hat{t}}_{i})/T)}\}}.$ (6)
@@
Diversity regularization. As local negative prompts are randomly initialized, more regularization has to be imposed on them to ensure their diversity. Therefore, in addition to the proposed contrastive regularization, we apply a diversity regularization on local negative prompts, as follows:
$\mathcal{L}_{\mathrm{reg}}=\frac{1}{N_{\mathrm{neg}}(N_{\mathrm{neg}}-1)/2}\sum_{1\leq i<j\leq N_{\mathrm{neg}}}\mathrm{sim}(\bm{\hat{t}}_{i},\bm{\hat{t}}_{j}),$ (7)
@@ 4.3 Regional OOD score
$S_{\mathrm{R-MCM}}(\bm{x})=S_{\mathrm{MCM}}(\bm{x})+\mathcal{T}_{k}^{\mathrm{mean}}\{\frac{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)}{\sum_{j=1}^{C}{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{l},\bm{t}_{j}})/T)+\sum_{j=1}^{N_{\mathrm{neg}}}{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{l},\bm{\hat{t}}_{j}})/T)}\},$ (9)
@@
$h_{i}^{\mathrm{g}}(\bm{x})=\mathrm{sim}(\bm{z}^{\mathrm{g}},\bm{t}_{i}),$ (10)
$h_{i}^{\mathrm{l}}(\bm{x})=\mathcal{T}_{k}^{\mathrm{mean}}\{\mathrm{exp}(sim(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\},$ (11)
@@
MSP 54.05 87.43 73.37 78.03 72.98 78.03 68.85 79.06 67.31 80.64
Energy 29.75 94.68 53.18 87.33 56.40 85.60 51.35 88.00 47.67 88.90
KNN 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
Textual-OE 29.61 94.74 57.12 87.34 66.82 83.71 79.29 77.76 58.21 85.88
@@
CoOp 18.95 95.52 29.58 92.90 38.72 89.64 48.03 85.87 33.82 90.98
Local-Prompt 9.65 97.87 20.40 95.57 29.39 92.67 51.20 88.00 27.66 93.53
LoCoOp 21.67 95.69 22.98 95.07 31.41 92.10 49.79 87.85 31.46 92.68
Local-Prompt+LoCoOp 12.81 97.29 19.34 95.85 27.53 92.97 45.51 89.99 26.29 94.03
@@
CoOp 14.60 96.62 28.48 92.65 36.49 89.98 43.13 88.03 30.67 91.82
Local-Prompt 8.71 98.10 23.97 94.85 32.50 92.32 47.93 89.04 28.27 93.58
LoCoOp 16.05 96.86 23.44 95.07 32.87 91.98 42.28 90.19 28.66 93.52
Local-Prompt+LoCoOp 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
5 Experiments
Datasets. Following existing works (Ming et al., 2022; Miyai et al., 2023b; a), large scale ImageNet-1K (Deng et al., 2009) along with a 100-category subset of it denoted as ImageNet-100 are used as ID dataset and OOD dataset is a combination of iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017) and Texture (Cimpoi et al., 2014). In addition, We also use two semantically similar subsets of ImageNet-1k, i.e., ImageNet-10 and ImageNet-20, to evaluate near OOD detection performance (Ming et al., 2022). Detailed information is provided in Appendix A.
Implementation details. We use CLIP-Base/16 as the backbone. Image encoder, text encoder, and global prompts are frozen and only local prompts are learnable. We set one local prompt for each known category (1000 in total for ImageNet-1k). $\lambda_{\mathrm{neg}}$ and $\lambda_{\mathrm{reg}}$ are $5$ and $0.5$, respectively. As the stability of temperature $T$ has been verified by previous research (Ming et al., 2022; Miyai et al., 2023b), we set $T$ to be 1 by default. More details are shown in Appendix B.
Evaluation metrics. We report the following metrics for evaluation: (1) the area under the receiver operating characteristic curve (AUROC); (2) false positive rate of OOD samples when true positive rate of ID samples is 95% (FPR95); (3) in-distribution data classification accuracy (ID accuracy).
5.1 Main results
ImageNet-1k as ID dataset. We report 4 and 16-shot results on four OOD datasets using ImageNet-1k as ID dataset. It can be seen in Tab. 1 that tuning local prompts only using hand-crafted global prompts can achieve competitive results, especially in datasets with challenging fine local textures like iNaturalist. Concretely, we get impressive progress in 4-shot tuning (outperforming by 3.80% on FPR95). In 16-shot setting, our approach gets competitive results as well. All results above strongly showcase the potential of regional enhancement for OOD detection as an orthogonal direction to global prompt optimization methods. Results of more shots are shown in Appendix C.
ID accuracy. In addition to OOD performance, it is also meaningful to evaluate ID accuracy as it is expected to ensure classification accuracy and separate outliers at the same time. However, weighing too much about separating outliers may narrow the inter-class distance, thus harming accuracy. Consequently, we evaluate ID accuracy for ImageNet-1k and the results shown in Tab. 3 conclude that our method significantly surpasses previous methods (1.02% and 2.38%, respectively). It demonstrates that the proposed regional enhancement is also beneficial to improve OOD separation while increasing inter-class distance, which is of great significance for real application.
Extensive to global prompts. Benefiting from the convenient integration with well-trained global prompts, we combine the advantage of trained global prompts from LoCoOp (Miyai et al., 2023a) without further design (Local-Prompt+LoCoOp) and get further substantial improvements on all datasets. For example, we surpass the state-of-the-art method by a large margin (5.17% on FPR95 and 1.35% on AUROC), even against previous methods with 16 shots. Specifically, extending local prompts to global prompts is helpful to datasets with major global representations like Texture, verifying the compatibility and potential of local prompt optimization. It is notable that the extension process is seamless and can be integrated with any global optimization prompts, strongly showing the extensibility of our approach. Density map in Fig. 3 also illustrates the superiority of our method. More visualization can be found in Appendix D.
@@ 5.1 Main results
MCM zero-shot 67.01
GL-MCM 67.01
KNN full-tuning 79.64
NPOS 79.42
LoCoOp 4-shot 69.50
Local-Prompt 70.52
LoCoOp 16-shot 71.86
Local-Prompt 74.24
Table 3: Comparison results of near OOD detection tasks. We report 16-shot results for few-shot methods. Our method can achieve performance against other methods.
@@
MCM 5.00 98.71 12.51 97.70 8.75 98.21
GL-MCM 10.10 98.04 9.00 98.62 9.55 98.33
LoCoOp 11.20 97.49 12.00 97.79 11.60 97.64
Local-Prompt 3.90 99.06 6.20 98.84 5.05 98.95
Table 4: Experiments on ImageNet-100. 4-shot results are reported for few-shot tuning.
@@
MCM 18.13 96.77 36.45 94.54 34.52 94.36 41.22 92.25 32.58 94.48
LoCoOp 44.94 93.31 30.70 94.70 34.31 93.93 55.92 90.01 41.47 92.99
Local-Prompt 6.99 98.05 24.34 96.03 26.04 95.49 36.32 93.67 23.42 95.81
Near OOD detection tasks. Near OOD task typically evaluates semantically similar categories and is relatively more challenging. Following existing work (Ming et al., 2022), we conduct experiments on ImageNet-10 vs. ImageNet-20 to verify the effectiveness of the proposed approach. Results in Tab. 3 indicate that our method exceeds previous methods by a large margin. It can be observed that previous few-shot method fails in near OOD tasks. We attribute it to the refinement of local prompts by local outlier knowledge, which greatly assists in distinguishing semantically similar categories. By contrast, relying solely on global prompts may confuse the distribution. Therefore, it underlines that our local outlier enhancement approach is especially effective in semantically near OOD tasks.
ImageNet-100 as ID dataset. We also conduct experiments using ImageNet-100 as ID dataset. We compare with other zero-shot and few-shot methods and results are reported in Tab. 4. The substantial performance improvements (average 9.16% reduction on FPR95 and 1.33% promotion on AUROC) strongly demonstrate the effectiveness and transferability of our approach.
Table 5: Comparison with NegLabel, which leverages powerful outlier samples.
Method iNaturalist SUN Places Texture Average
[…]
NegLabel (w outlier) 1.91 99.49 20.53 95.49 35.59 91.64 43.56 90.22 25.40 94.21
Local-Prompt (w/o outlier) 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
Comparison with outlier methods. NegLabel (Jiang et al., 2024) designs a novel scheme for OOD score with negative labels. It leverages real outlier information with negative labels from extensive corpus databases. This kind of knowledge helps to a great extent pointed out by OE (Hendrycks et al., 2018) and is inconsistent with real-world application, where negative categories are infinite. As can be seen from Tab. 5, we observe that (1) although without outlier, our model achieves better average performance against NegLabel; (2) our method achieves better balance between different kinds of OOD datasets, which strongly validates the effectiveness of incorporating local information and strengthens its application in diverse and infinite read-world scenarios.
@@
Figure 4: Ablation study of different OOD score strategies. LNP denotes local negative prompts.
5.2 Ablation Study
We conduct various and comprehensive ablation studies to investigate the effectiveness of each component. We train a 4-shot model to showcase the function of components and hyper-parameters unless otherwise stated. More studies about training time and coefficient are shown in Appendix C.
Effectiveness of loss components. We conduct experiments to analyze the effectiveness of each loss component and results in Tab. 6 imply that local prompts optimization (baseline) achieves comparable results with global prompts optimization methods, strongly demonstrating the effectiveness of our proposed regional enhancement strategy. Furthermore, taking hard negative samples together with local negative prompts $\bm{\hat{t}}$ to optimize OOD-considered representation is beneficial to the distinction of outliers. Outcome without regularization loss is better in FPR95 but slightly inferior in AUROC. Considering overall performance and the fact that the optimization of $\bm{\hat{t}}$ relies heavily on the initialization without regularization, we reserve regularization loss for stability and convergence.